The density of NO2 in a 2.50 L tank at 720.0 torr and 35.0°C is __________ g/L.
Answer #1

The equation to be used to solve the equations is the ideal gas equation:

PV = nRT where P = pressure(in atm), V = volume(in L) , n = number of moles of gas, R = Universal Gas Constant ( 0.0821 L atm mol-1 K-1) and T = Temperature (in K).

Now,

Density (d) = Mass(m)/Volume

That is , d = m/V

Also, mass ( in g) = Number of moles ( n) x Molar mass ( M)

So, m = nM. Putting in formula for density :

d = nM/V

or, ( n/V) = d/M

The ideal gas equation can be modified as:

PV = nRT

Or, n/V = P/RT

Or, d/M = P/RT

So, d = PM /RT

Given :

P = 720.0 torr = 0.9474 atm ( relation: 760 torr = 1 atm)

M = molar mass of NO2= 46 g/mol

T = 35.0 C = ( 35 + 273 ) K = 308 K

Putting the values :

Density (0.9474 atm) (46 g/mol) Latm (0.0821 (308 K) molk

Solving, density of NO2 = 1.723 g/L